OpenStudy (anonymous):

integrate tanx^4 sec x^3

OpenStudy (shamil98):

If you remember \[(secx)' = \tan(x)\sec(x)\] In this case you have a composite function (exponentials) instead of just a regular x.

OpenStudy (shamil98):

Wait, is this \[\int\limits_{}^{} \tan^4 x \sec^3x dx\] or \[\int\limits_{}^{} \tan(x^4) \sec(x^3) dx\]

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

I wish it was the second one

OpenStudy (anonymous):

forgot the ( )

OpenStudy (shamil98):

You can rewrite it as: \[\int\limits_{}^{} \tan^2(x)* \tan^2(x) * \sec^2 (x) * \sec(x) dx\] and since \[\sec^2(x) = 1 + \tan^2 (x)\] from here it's just u-sub..

OpenStudy (anonymous):

using identity right

OpenStudy (shamil98):

yes that was a trig identity.

OpenStudy (anonymous):

thx was just stump been a min I looked over this

OpenStudy (shamil98):

you're welcome :)